A monopolist’s demand and Average total has cost functions are given by

Q=250-5P

ATC =10 +40/Q

A. What will be the profit maximizing level of price and output?

B. What will be the firm’s total profit if it exercises first degree price discrimination?

C. Assume the monopoly can sell his product in two separated markets with

the following demand functions

Q1 =110-P1

Q2=140-4P2

A. Determine the profit maximizing level output in each market

B. What price should the monopoly sell each unit of output in each market?

C. In which market is the price high? Why?


1
Expert's answer
2021-12-22T13:22:40-0500

A)

TC= ACT*Q

=(10+40/Q)Q

=10Q+40

Q=250-5P

5P=250-Q

P=50-0.2Q

TR= P*Q

=(50-0.2Q)Q

=50-0.2Q2

MR= 50-0.4Q

MC=10

For profit maximization,

MR=MC

50-0.4Q=10

0.4Q=40

Q= 100

p=50-0.2(100)

= 30

B) Firms total profit

"\Pi" =TR-TC

TR=50(100)-0.2(1002)

=3000

TC=10(100)+40

=1040

"\Pi"=3000-1040

=1960

C)( a) profit Maximizing in each market

Q1=110-P1

P1=110-Q1

Q2=140-4P2

4P2=140-Q2

P2=35-0.25Q2

For profit Maximization in each marlket,

MR1=MC

MR2=MC

TR1=P1Q1

=(110-Q1)Q1=110Q1-Q12

110Q1-Q12

MR1=110-2Q1

110-2Q1=10

2Q1=100

Q1=50

MR2=MC

TR2=P1Q1

=(35-0.25Q2)Q2

=35Q2-0.25Q22

MR2=35-0.5Q2

35-0.5Q2=10

0.5Q2=25

Q2=50

B) What price the monopoly should charge in each market

P1=110-Q1

=110-50= 60

P2= 35-0.25Q2

=35-0.25(50)

= 22.5

C) The price is higher in market 1 because it is price inelastic